6) 17pt) Solve the following IVP. y +4y = f(t) y(O) = 0 ;(0) = 0 if t <3 = (exp(0.2t) if t23

wondering how to solve this using the Laplace transformation. thank you
Answer #1

If you have any doutbs in the solution please ask me in commments ...here i used laplace method to solve the ivp and crossed check it also

** To solve the following IVP, y + 4y = f1t) y(0) = 0, y(0) = 0 if t <3 fit) es ; exp 10.2 t) ; if to3 O NOW as we know thaso the given IVP will be, 775 y + uy e Ult-3) 4(0)-0, y0) = 0 Lly) = y(s). NOW By apply Laplace Both side assume L(4) + Uflt) = la (152+4) (5-5) А + (52+4)(5-5) Bst C (S- S?+ 4 As?+ 4A + Bs? + CS - see A+B=0, § -01, 6-= 1 ЧА - = By compare → B =so the solution of the given IVP; +15 25 e e+315 es 25 101 cos 2 (4-3) + 5 sin sin 2(t-3) ylt) = +-3). 202 101 Answer